# Formal Analysis of the Contract Automata Runtime Environment with Uppaal: Modelling, Verification and Testing

### v2

@@ 1. Introduction
Behavioural contracts [7] have been introduced to formally describe the interactions among services, to enable to reason formally about well-behaving properties of their composition. Examples of properties are agreement among the parties or reachability of target states.
Contract automata are a dialect of finite state automata used to specify behavioural contracts formally in terms of offers and requests [21]. A composition of contracts is in agreement when all requests are matched by corresponding offers of other contracts. A composition can be refined to one in agreement using the orchestration synthesis algorithm [17, 20], a variation of the synthesis algorithm from supervisory control theory [61]. The Contract Automata Runtime Environment (CARE[15] provides a middleware to coordinate the services implementing contracts. In CARE, each transition of the orchestration automaton is executed by a series of interactions among the orchestrator and the CARE services, implemented using Java TCP/IP sockets. These interactions may vary according to the specific configuration chosen among those provided by CARE. In [15] the algorithms implemented in CARE are proved to enforce the adherence of each contract specification to its CARE implementation (basically, the control flow of the application follows the synthesised orchestration automaton).
In this paper, we describe the modelling, verification and testing of the low-level interactions among CARE services and the orchestrator. This aspect is no less important, as witnessed by known cases of algorithms proved to be correct (e.g., the Byzantine distributed consensus [48]) and whose low-level communication implementations were found to have issues, e.g. deadlocks [62]. We verify several properties, including the absence of deadlocks, absence of undelivered messages, and reachability of target states.
@@
Finally, this paper tackles the challenge of providing a full-fledged model-based development and formal methods approach [37]. The final application has been graphically modelled at an abstract level, formally verified and tested using the formal model. The criteria followed for abstracting away irrelevant details are discussed together with the issues that have been found and fixed thanks to the formal modelling, verification and testing.
All models, formulas, and logs are publicly available in [10], together with traceability and model-based testing information connecting the model to the source code.
This work is an extension of the conference paper [11]. The extension includes the addition of more figures and tables, as well as further explanations throughout the paper. The most significant addition concerns the testing phase, which was only briefly mentioned in [11]. Testing is now thoroughly described in Section 7. We cover both the generation of abstract tests from the Uppaal model and the process of concretising these tests in the real application.
@@ 2. Related work
Several applications of Uppaal to various case studies are available in the literature, including land transport [19], maritime transport [64], medical systems [52], and autonomous agents path planning [39]. These case studies, along with the present paper, adopt a model-based approach, wherein partial representations of the applications are created using models.
@@
A recent survey conducted on formal methods [36] reveals that numerous publications lack a direct correlation between the concrete implementation and its abstract model. This holds especially when the systems are only envisioned, yet to be realised, such as for example the ERTMS L3 railway signalling system [8, 18]. This is also the case for industrial systems whose implementations are non-disclosed [43]. In all these cases, determining the accuracy of the model in relation to the actual system and assessing the appropriateness of the chosen level of abstraction becomes challenging. Furthermore, measuring the influence of the formal modelling and verification phase on the analyzed system poses difficulties.
@@
Indeed, a recent survey on formal methods in the transport domain [36] identifies the need “to lower the degree of abstraction, by applying formal methods and tools to development phases that are closer to software development”. This is also confirmed by a recent survey on formal methods among 130 high-profile experts [37], where nearly half of them believe that the most likely future users of formal methods are “a small number of skilled experts” as long as “skills like modelling, specification, abstraction are involved”.
@@
In contrast to the previously mentioned literature, in this paper we present a bottom-up formal analysis of an established open-source system that has already been developed [15]. The availability of the source code further enables us to establish a connection between the abstract formal model and the actual source code. This capability facilitates the precise identification of specific aspects of the real system that have been abstracted in the formal model. Additionally, it allows us to validate the appropriateness of the chosen level of abstraction.
@@
Recently, the offline model-based testing functionality of Uppaal, called Yggdrasil [41], has been used in [46] for testing an industrial real-time automotive turn indicator system, and in [40] for testing EIP-1559 Ethereum smart contracts. The usage of the Uppaal suite, and more specifically Yggdrasil, in three industrial case studies (i.e., medical devices, an automotive system and a pump manufacturing) is reviewed in [50]. Many of the above applications of model-based testing using Uppaal concern industrial, non-disclosed case studies. Therefore, only the models and abstract test cases are reported, whilst concrete test cases and source code have not been made public. In comparison, CARE is an open-source application. The concrete tests generated from the Uppaal models discussed in this paper, as well as the system under testing, are available at [10]. To the best of our knowledge, there are no other non-trivial open-source applications for which their Uppaal formal model is openly accessible and directly connected to the source code through traceability and model-based testing. This contribution assists in linking formal methods, particularly Uppaal, to the software development process.
@@
The contract automata approach is closer to [47], where behavioural types are expressed as finite state automata of Mungo, called typestates [65, 4]. Similarly to CARE, in Mungo finite state automata are used as behaviour assigned to Java classes (one automaton per class), with transition labels corresponding to methods of the classes. A tool to translate typestates into automata was presented in [66]. CATApp is a graphical front-end tool for designing contract automata [9]. A tool similar to Mungo is JaTyC (Java Typestate Checker) [6].
@@
Other model-based verification approaches for service-oriented systems and distributed middleware are surveyed in [60, 31]. In the literature, there exist many formalisms for modelling and analysing the behaviour of services, ranging from behavioural type systems, including behavioural contracts [29, 1, 49] and session types [26, 42, 34, 28, 55], to automata-based formalisms, including interface automata [3] and (timed) (I/O) automata [54, 5, 33]. Foundational models for service contracts and session types are surveyed in [23, 7, 44]. Sessions and session types [26, 42, 34, 28, 55] have been introduced to reason over the behaviour of services in terms of their interactions. Compared to contract automata, behavioural contracts and session types use process-algebraic frameworks, instead of finite state automata. Indeed, contract automata are similar to other software engineering-based formalisms such as I/O automata [54, 5, 33].
@@
In the literature there are various attempts to model applications that, similarly to CARE, communicates through TCP/IP sockets, also using Uppaal [67, 35, 63, 68]. In [30] the verification of the transport layer of a TCP/IP protocol stack implementation in C is performed with a translation into SPARK, where contracts are modelling the corresponding behaviour. A method to obtain SPIN formal models from TCP/IP socket applications implemented in C is presented in [27] where the C application is embedded into a PROMELA model equipped with a formalisation of the underlying TCP/IP socket API. Similarly, we provide a model of the underlying socket communications but we also modelled aspects of the application that are targeted by our analysis whilst unnecessary details are abstracted away. Another approach [53] presents an automatic translation from MAUDE specifications of distributed applications to distributed MAUDE implementations using TCP/IP sockets, where the implementations are formally proven to be in correspondence with the specifications. Our approach is bottom-up and is based on Uppaal modelling of an already realised application.
@@
As reported in [25], the question of qualification and validation of formal methods tools is “absolutely crucial”. Other tools for behavioural contracts are present in the literature [38, 57, 58], but differently from CARE, many of these implementations have not undergone a process of formal verification.
@@ 3.1. Contract Automata
In a composition of contracts various properties can be analysed [21]. For example, the property of agreement requires to match all request actions, whereas offer actions can remain unmatched. The synthesis of the orchestration in agreement produces a sub-automaton of the composition where all services can match their requests with corresponding offers to reach a final state. Thus, in the orchestration in agreement labels of transitions are only matches or offers [20]. The contract automaton in Figure 11 right is an orchestration in agreement.
Contract automata and their functionalities are implemented in a software artefact, called Contract Automata Library (CATLib[13]. With CATLib it is possible to specify, compose, and synthesise specifications given as contract automata. Indeed, CARE uses the facilities offered by CATLib for, among other things, composing contracts and synthesising orchestrations. This software artefact is a by-product of our scientific research on behavioural contracts and implements results that have previously been formally specified in several publications (cf., e.g., [17, 20, 21]). Scalability features offered by CATLib include a bounded on-the-fly state-space generation optimised with pruning of redundant transitions and parallel streams computations [14, 16]. The software is open source [13], it has been developed using principles of model-based software engineering [12] and it has been extensively validated using various testing and analysis tools to increase the confidence on the reliability of the library [13].
@@ 3.2. Uppaal
Uppaal SMC [32] is a variant of Uppaal [24], which is a well-known toolbox for the verification of real-time systems. Uppaal models are stochastic timed automata, in which non-determinism is replaced with probabilistic choices, and time delays with probability distributions (uniform for bounded time and exponential for unbounded time). These automata may communicate only via broadcast channels and shared variables. In this paper, we will use both exhaustive and statistical model checking. Statistical Model Checking (SMC) [51] involves running a sufficient number of (probabilistic) simulations of a system model to obtain statistical evidence (with a predefined level of statistical confidence) of the quantitative properties to be checked. Monte Carlo estimation with Chernoff-Hoeffding bound executes $N=\lceil(ln(2)-ln(\alpha))/(2\varepsilon^{2})\rceil$ simulations $\rho_{i}$, $i\!\in\!1...N$, to provide the interval $[p^{\prime}-\varepsilon,p^{\prime}+\varepsilon]$ with confidence $1-\alpha$. Here, $p^{\prime}=(\#\{\rho_{i}\!\mid\!\rho_{i}\!\models\!\varphi\})/N$, i.e., $\mathtt{Pr}(|p^{\prime}-p|\leq\varepsilon)\geq 1-\alpha$ where $p$ is the unknown value of the formula $\varphi$ being estimated statistically and $\varepsilon$ and $\alpha$ are the user-defined precision and confidence, respectively. Crucially, the number of simulations used to estimate a formula is independent of the model’s size and depends only on the parameters $\alpha$ and $\varepsilon$. In practice the number of simulations required by Uppaal to reach a specific confidence level is optimized and is thus lower than the above theoretical bound. Uppaal supports template automata used to instantiate different copies (in different experiments) of the same automaton, distinguishable by their parameters. The selection of Uppaal as the chosen tool is influenced by several factors. These include its extensive adoption by the community, expertise of the author, usability, primitive support for real-time and stochastic modelling, probabilistic and non-deterministic choices, and capabilities for statistical model checking, simulation, and model-based testing.
@@ 3.3. CARE
The Contract Automata Runtime Environment (CARE[15] provides facilities for pairing the contract automata specifications with actual implementations of service-based applications. Note that our purpose is not to formally analyse the applications created using CARE, but to formally analyse CARE itself. The formal verification of applications developed using CARE is a consequence of the reliance of CARE on the formal guarantees provided by contract automata [15], along with the formal verification of CARE as an application. This paper addresses the latter aspect. This software is organised into classes for the orchestrated services (cf. Figure 2) and classes for the orchestrator (cf. Figure 3). The two core abstract Java classes are the RunnableOrchestration and the RunnableOrchestratedContract.
@@
In Figure 2, RunnableOrchestratedContract is an abstract class that implements an executable wrapper responsible for pairing its contract automaton specification (instance variable contract storing a contract automaton) with an implementation provided as a Java class (instance variable service implementing the service), where each action of the automaton contract is in correspondence with a method of the service class. The types of the methods of a matching offer and request must also be in correspondence. Each time a new orchestration involving the service is initiated, the RunnableOrchestratedContract creates a new service. This service remains in a waiting state to receive commands from the orchestrator, which then triggers the execution of the corresponding methods. In case the orchestrator requires to perform an action not prescribed by its service contract, then a ContractViolationException will be raised by the service. As demonstrated in [15], this exception will never be raised if the orchestration in agreement is synthesised using the algorithms of contract automata.
@@
Two aspects to implement for both the orchestrator and the services are choices and termination (through the abstract method choice). Indeed, during the execution of an orchestration, the selection of the transition to execute in the presence of multiple enabled transitions is implemented through this abstract method. CARE is equipped with default implementations, but can be extended (by implementing the relative interfaces and abstract methods) to include other options, other than the default ones. Currently, a so-called ‘dictatorial’ choice (i.e., an internal choice of the orchestrator, external for the services) and a so-called ‘majoritarian’ choice (services vote and the majority wins) are two implemented options. For the services, MajoritarianChoiceRunnableOrchestratedContract and DictatorialChoiceRunnableOrchestratedContract are the two classes specialising RunnableOrchestratedContract according to how the choice is handled and implementing the abstract methods, whilst MajoritarianChoiceRunnableOrchestration and DictatorialChoiceRunnableOrchestration are those specialising RunnableOrchestration. CARE also provides default implementations for the low-level message exchanges. Currently, the two available options are the ‘centralised’ action, where the orchestrator acts as a proxy, and the ‘distributed’ action, where two services matching their actions directly interact with each other once the orchestrator has made them aware of a matching partner and its address/port. Accordingly, each RunnableOrchestratedContract has an OrchestratedAction, and RunnableOrchestration has an OrchestratorAction (instance variable act) used to implement the corresponding actions that can be either distributed or centralised according to the current implementation.
@@
We discuss an example of interactions between two services and an orchestrator in CARE. The example comprehends both the CentralisedAction and DistributedAction implementations of CARE. The sequence of service invocations is displayed in Figure 4. In this example, the orchestrator is executing a transition of the orchestration automaton that is labelled with the match [?coffee,!coffee], in which Alice is requesting a coffee and Bob is offering a coffee. As stated above, the RunnableOrchestratedContract class of both Alice and Bob has two instance variables (contract and service, resp.) containing the contract automaton object and the implementation provided as a Java class. In this example, the implementation class of Alice has a method Integer coffee(String arg) that is paired with the corresponding request action ?coffee of her contract. Similarly, the implementation class of Bob has a method String coffee(Integer arg) that is paired with the corresponding offer action !coffee of his contract.

### v3

@@ 1. Introduction
Behavioural contracts [BCZ15] have been introduced to formally describe the interactions among services, to enable to reason formally about well-behaving properties of their composition. Examples of properties are agreement among the parties or reachability of target states.
Contract automata are a dialect of finite state automata used to specify behavioural contracts formally in terms of offers and requests [BDF16]. A composition of contracts is in agreement when all requests are matched by corresponding offers of other contracts. A composition can be refined to one in agreement using the orchestration synthesis algorithm [BtBD+20, BtBP20], a variation of the synthesis algorithm from supervisory control theory [RW87]. The Contract Automata Runtime Environment (CARE[BtB23a] provides a middleware to coordinate the services implementing contracts. In CARE, each transition of the orchestration automaton is executed by a series of interactions among the orchestrator and the CARE services, implemented using Java TCP/IP sockets. These interactions may vary according to the specific configuration chosen among those provided by CARE. In [BtB23a] the algorithms implemented in CARE are proved to enforce the adherence of each contract specification to its CARE implementation (basically, the control flow of the application follows the synthesised orchestration automaton).
In this paper, we describe the modelling, verification and testing of the low-level interactions among CARE services and the orchestrator. This aspect is no less important, as witnessed by known cases of algorithms proved to be correct (e.g., the Byzantine distributed consensus [LSP82]) and whose low-level communication implementations were found to have issues, e.g. deadlocks [RCS+22]. We verify several properties, including the absence of deadlocks, absence of undelivered messages, and reachability of target states.
@@
Finally, this paper tackles the challenge of providing a full-fledged model-based development and formal methods approach [GtBvdP20]. The final application has been graphically modelled at an abstract level, formally verified and tested using the formal model. The criteria followed for abstracting away irrelevant details are discussed together with the issues that have been found and fixed thanks to the formal modelling, verification and testing.
All models, formulas, and logs are publicly available in [Basb], together with traceability and model-based testing information connecting the model to the source code.
This work is an extension of the conference paper [Bas24]. The extension includes the addition of more figures and tables, as well as further explanations throughout the paper. The most significant addition concerns the testing phase, which was only briefly mentioned in [Bas24]. Testing is now thoroughly described in Section 7. We cover both the generation of abstract tests from the Uppaal model and the process of concretising these tests in the real application.
@@ 2. Related work
Several applications of Uppaal to various case studies are available in the literature, including land transport [BtBL20], maritime transport [SVW20], medical systems [LRN+22], and autonomous agents path planning [GJP+22]. These case studies, along with the present paper, adopt a model-based approach, wherein partial representations of the applications are created using models.
@@
A recent survey conducted on formal methods [FtB22] reveals that numerous publications lack a direct correlation between the concrete implementation and its abstract model. This holds especially when the systems are only envisioned, yet to be realised, such as for example the ERTMS L3 railway signalling system [BtBC18, BtBFL22]. This is also the case for industrial systems whose implementations are non-disclosed [HMG+23]. In all these cases, determining the accuracy of the model in relation to the actual system and assessing the appropriateness of the chosen level of abstraction becomes challenging. Furthermore, measuring the influence of the formal modelling and verification phase on the analyzed system poses difficulties.
@@
Indeed, a recent survey on formal methods in the transport domain [FtB22] identifies the need “to lower the degree of abstraction, by applying formal methods and tools to development phases that are closer to software development”. This is also confirmed by a recent survey on formal methods among 130 high-profile experts [GtBvdP20], where nearly half of them believe that the most likely future users of formal methods are “a small number of skilled experts” as long as “skills like modelling, specification, abstraction are involved”.
@@
In contrast to the previously mentioned literature, in this paper we present a bottom-up formal analysis of an established open-source system that has already been developed [BtB23a]. The availability of the source code further enables us to establish a connection between the abstract formal model and the actual source code. This capability facilitates the precise identification of specific aspects of the real system that have been abstracted in the formal model. Additionally, it allows us to validate the appropriateness of the chosen level of abstraction.
@@
Recently, Yggdrasil [HLM+08], the offline model-based testing functionality of Uppaal, has been used in [KLN+15] for testing an industrial real-time automotive turn indicator system, and in [HL23] for testing EIP-1559 Ethereum smart contracts. The usage of the Uppaal suite, and more specifically Yggdrasil, in three industrial case studies (i.e., medical devices, an automotive system and a pump manufacturing) is reviewed in [LLN18]. Many of the above applications of model-based testing using Uppaal concern industrial, non-disclosed case studies. Therefore, only the models and abstract test cases are reported, whilst concrete test cases and source code have not been made public. In comparison, CARE is an open-source application. The concrete tests generated from the Uppaal models discussed in this paper, as well as the system under testing, are available at [Basb]. To the best of our knowledge, there are no other non-trivial open-source applications for which their Uppaal formal model is openly accessible and directly connected to the source code through traceability and model-based testing. This contribution assists in linking formal methods, particularly Uppaal, to the software development process.
@@
The contract automata approach is closer to [KDPG18], where behavioural types are expressed as finite state automata of Mungo, called typestates [SY86, AD24]. Similarly to CARE, in Mungo finite state automata are used as behaviour assigned to Java classes (one automaton per class), with transition labels corresponding to methods of the classes. A tool to translate typestates into automata was presented in [TMR20]. CATApp is a graphical front-end tool for designing contract automata [Basa]. A tool similar to Mungo is JaTyC (Java Typestate Checker) [BBG+22].
@@
Other model-based verification approaches for service-oriented systems and distributed middleware are surveyed in [RG21, CHLR23]. In the literature, there exist many formalisms for modelling and analysing the behaviour of services, ranging from behavioural type systems, including behavioural contracts [CGP09, ABZ13, LP15] and session types [BLMT08, HYC08, DCd10, CDCP12, MNF13], to automata-based formalisms, including interface automata [dAH01] and (timed) (I/O) automata [LT89, AD94, DLL+10]. Foundational models for service contracts and session types are surveyed in [tBBG07, BCZ15, HLV+16]. Sessions and session types [BLMT08, HYC08, DCd10, CDCP12, MNF13] have been introduced to reason over the behaviour of services in terms of their interactions. Compared to contract automata, behavioural contracts and session types use process-algebraic frameworks, instead of finite state automata. Indeed, contract automata are similar to other software engineering-based formalisms such as I/O automata [LT89, AD94, DLL+10].
@@
In the literature there are various attempts to model applications that, similarly to CARE, communicates through TCP/IP sockets, also using Uppaal [WWH+22, FZL18, SF17, WVHFM06]. In [CGMZ21] the verification of the transport layer of a TCP/IP protocol stack implementation in C is performed with a translation into SPARK, where contracts are modelling the corresponding behaviour. A method to obtain SPIN formal models from TCP/IP socket applications implemented in C is presented in [dlCGMS09] where the C application is embedded into a PROMELA model equipped with a formalisation of the underlying TCP/IP socket API. Similarly, we provide a model of the underlying socket communications but we also modelled aspects of the application that are targeted by our analysis whilst unnecessary details are abstracted away. Another approach [LSM+20] presents an automatic translation from MAUDE specifications of distributed applications to distributed MAUDE implementations using TCP/IP sockets, where the implementations are formally proven to be in correspondence with the specifications. Our approach is bottom-up and is based on Uppaal modelling of an already realised application.
@@
As reported in [Bou15], the question of qualification and validation of formal methods tools is “absolutely crucial”. Other tools for behavioural contracts are present in the literature [GR17, OPB+21, PMT24], but differently from CARE, many of these implementations have not undergone a process of formal verification.
@@ 3.1. Contract Automata
In a composition of contracts various properties can be analysed [BDF16]. For example, the property of agreement requires to match all request actions, whereas offer actions can remain unmatched. The synthesis of the orchestration in agreement produces a sub-automaton of the composition where all services can match their requests with corresponding offers to reach a final state. Thus, in the orchestration in agreement labels of transitions are only matches or offers [BtBP20]. The contract automaton in Figure 11 right is an orchestration in agreement.
Contract automata and their functionalities are implemented in a software artefact, called Contract Automata Library (CATLib[BtB22]. With CATLib it is possible to specify, compose, and synthesise specifications given as contract automata. Indeed, CARE uses the facilities offered by CATLib for, among other things, composing contracts and synthesising orchestrations. This software artefact is a by-product of our scientific research on behavioural contracts and implements results that have previously been formally specified in several publications (cf., e.g., [BtBD+20, BtBP20, BDF16]). Scalability features offered by CATLib include a bounded on-the-fly state-space generation optimised with pruning of redundant transitions and parallel streams computations [BtB23b, BtB24]. The software is open source [BtB22], it has been developed using principles of model-based software engineering [BtB21] and it has been extensively validated using various testing and analysis tools to increase the confidence on the reliability of the library [BtB22].
@@ 3.2. Uppaal
Uppaal SMC [DLL+15] is a variant of Uppaal [BDL+06], which is a well-known toolbox for the verification of real-time systems. Uppaal models are stochastic timed automata, in which non-determinism is replaced with probabilistic choices, and time delays with probability distributions (uniform for bounded time and exponential for unbounded time). These automata may communicate only via broadcast channels and shared variables. In this paper, we will use both exhaustive and statistical model checking. Statistical Model Checking (SMC) [LLT+19] involves running a sufficient number of (probabilistic) simulations of a system model to obtain statistical evidence (with a predefined level of statistical confidence) of the quantitative properties to be checked. Monte Carlo estimation with Chernoff-Hoeffding bound executes $N=\lceil(ln(2)-ln(\alpha))/(2\varepsilon^{2})\rceil$ simulations $\rho_{i}$, $i\!\in\!1...N$, to provide the interval $p^{\prime}-\varepsilon,p^{\prime}+\varepsilon]$ with confidence $1-\alpha$. Here, $p^{\prime}=(\#\{\rho_{i}\!\mid\!\rho_{i}\!\models\!\varphi\})/N$, i.e., $\mathtt{Pr}(|p^{\prime}-p|\leq\varepsilon)\geq 1-\alpha$ where $p$ is the unknown value of the formula $\varphi$ being estimated statistically and $\varepsilon$ and $\alpha$ are the user-defined precision and confidence, respectively. Crucially, the number of simulations used to estimate a formula is independent of the model’s size and depends only on the parameters $\alpha$ and $\varepsilon$. In practice the number of simulations required by Uppaal to reach a specific confidence level is optimized and is thus lower than the above theoretical bound. Uppaal supports template automata used to instantiate different copies (in different experiments) of the same automaton, distinguishable by their parameters. The selection of Uppaal as the chosen tool is influenced by several factors. These include its extensive adoption by the community, expertise of the author, usability, primitive support for real-time and stochastic modelling, probabilistic and non-deterministic choices, and capabilities for statistical model checking, simulation, and model-based testing.
@@ 3.3. CARE
The Contract Automata Runtime Environment (CARE[BtB23a] provides facilities for pairing the contract automata specifications with actual implementations of service-based applications. Note that our purpose is not to formally analyse the applications created using CARE, but to formally analyse CARE itself. The formal verification of applications developed using CARE is a consequence of the reliance of CARE on the formal guarantees provided by contract automata [BtB23a], along with the formal verification of CARE as an application. This paper addresses the latter aspect. This software is organised into classes for the orchestrated services (cf. Figure 2) and classes for the orchestrator (cf. Figure 3). The two core abstract Java classes are the RunnableOrchestration and the RunnableOrchestratedContract.
@@
In Figure 2, RunnableOrchestratedContract is an abstract class that implements an executable wrapper responsible for pairing its contract automaton specification (instance variable contract storing a contract automaton) with an implementation provided as a Java class (instance variable service implementing the service), where each action of the automaton contract is in correspondence with a method of the service class. The types of the methods of a matching offer and request must also be in correspondence. Each time a new orchestration involving the service is initiated, the RunnableOrchestratedContract creates a new service. This service remains in a waiting state to receive commands from the orchestrator, which then triggers the execution of the corresponding methods. In case the orchestrator requires to perform an action not prescribed by its service contract, then a ContractViolationException will be raised by the service. As demonstrated in [BtB23a], this exception will never be raised if the orchestration in agreement is synthesised using the algorithms of contract automata.
@@
Two aspects to implement for both the orchestrator and the services are choices and termination (through the abstract method choice). Indeed, during the execution of an orchestration, the selection of the transition to execute in the presence of multiple enabled transitions is implemented through this abstract method. CARE is equipped with default implementations, but can be extended (by implementing the relative interfaces and abstract methods) to include other options, other than the default ones. Currently, a so-called ‘dictatorial’ choice (i.e., an internal choice of the orchestrator, external for the services) and a so-called ‘majoritarian’ choice (services vote and the majority wins) are two implemented options. For the services, MajoritarianChoiceRunnableOrchestratedContract and DictatorialChoiceRunnableOrchestratedContract are the two classes specialising RunnableOrchestratedContract according to how the choice is handled and implementing the abstract methods, whilst MajoritarianChoiceRunnableOrchestration and DictatorialChoiceRunnableOrchestration are specialising RunnableOrchestration.
CARE also provides default implementations for the low-level message exchanges. Currently, the two available options are the ‘centralised’ action, where the orchestrator acts as a proxy, and the ‘distributed’ action, where two services matching their actions directly interact with each other once the orchestrator has made them aware of a matching partner and its address/port. Accordingly, each RunnableOrchestratedContract has an OrchestratedAction, and RunnableOrchestration has an OrchestratorAction (instance variable act) used to implement the corresponding actions that can be either distributed or centralised according to the current implementation.
@@
We discuss an example of interactions between two services and an orchestrator. The example comprehends both the CentralisedAction and DistributedAction implementations of CARE. The sequence of service invocations is displayed in Figure 4. In this example, the orchestrator is executing a transition of the orchestration automaton that is labelled with the match [?coffee,!coffee], in which Alice is requesting a coffee and Bob is offering a coffee. As stated above, the RunnableOrchestratedContract class of both Alice and Bob has two instance variables (contract and service, resp.) containing the contract automaton object and the implementation provided as a Java class. In this example, the implementation class of Alice has a method Integer coffee(String arg) that is paired with the corresponding request action ?coffee of her contract. Similarly, the implementation class of Bob has a method String coffee(Integer arg) that is paired with the corresponding offer action !coffee of his contract.